The average speed of a tractor is two-fifths the average speed of a car .The car covers 450 kms, in 6 hours. How much distance will the tractor cover in 8 hours ?
Answer by lwsshak3(11628)

The average speed of a tractor is two-fifths the average speed of a car .The car covers 450 kms, in 6 hours. How much distance will the tractor cover in 8 hours ?
***
speed=distance/travel time
speed of car=450/6=75 km/hr
speed of tractor=(2/5)*speed of car=(2/5)*75=30 km/hr
distance tractor will cover in 8 hours=30*8=240 km